OpenStudy (poetryprincess):

What is 15 1/3 in radical form?

Nnesha (nnesha):

\[\huge\rm 15^\frac{ 1 }{ 3 }\] ?? like this ??

OpenStudy (poetryprincess):

Yes and the answer choices are: \[A. \sqrt[15]{3}\] \[B. \sqrt[1/3]{15}\] \[C. \sqrt[3]{15}\]

Nnesha (nnesha):

hint: \[\huge\rm \sqrt[n]{x^m}= x^\frac{ m }{ n }\]

OpenStudy (anonymous):

so it would be 3 sqrrt 15 to the power of 1 = 15 1/3

Nnesha (nnesha):

you can change exponent to root radical form) so how would you write \[\huge\rm 15^\frac{ 1 }{ 3 }\] ??

OpenStudy (anonymous):

the answer is 46/3 or 15.333333333333333 and

OpenStudy (anonymous):

is it correct

OpenStudy (poetryprincess):

C is the correct answer.